# A-Bench: Are LMMs Masters at Evaluating AI-generated Images?

### v1

@@ 4.2 Findings of A-Bench
IDEFICS-2 (Mistral-7B-Instruct-v0.2) 61.18% 68.86% 47.36% 59.00%
DeepSeek-VL-7B 55.89% 53.81% 47.56% 52.40%
InternLM-XComposer2-VL (InternLM2) 62.24% 63.32% 50.31% 58.55%
LLaVA-NeXT (Llama3-8B) 58.54% 48.52% 52.03% 53.09%
LLaVA-NeXT (Qwen-72B) 59.96% 55.51% 59.76% 58.45%
LLaVA-NeXT (Qwen-110B) 64.63% 57.20% 63.62% 61.88%
mPLUG-Owl2 (LLaMA-7B) 57.93% 54.45% 53.86% 55.43%
LLaVA-v1.5 (Vicuna-v1.5-7B) 45.93% 41.31% 54.47% 47.32%
LLaVA-v1.5 (Vicuna-v1.5-13B) 46.14% 41.31% 48.17% 45.26%
CogVLM-17B (Vicuna-v1.5-7B) 54.88% 48.31% 52.44% 51.92%
Qwen-VL (Qwen-7B) 49.59% 34.32% 50.41% 44.92%
BakLLava (Mistral-7B) 47.84% 33.33% 48.44% 43.37%
Fuyu-8B (Persimmon-8B) 44.69% 30.28% 45.67% 40.27%
random guess 31.89% 32.91% 33.11% 32.62%
@@ A.1 AIGIs Collection
    Bag-of-Words -> Attributes Awareness: Six logs in a woodpile, stacked so tightly that they seem to form a solid block.
@@
Afterward, we use the collected prompts to create AIGIs. 15 text-to-image generation models are selected, which include: Dreamlike [dreamlike art, 2023], Pixart $\alpha$ Chen et al. [2023b], Playground v2 PlaygroundAI [2023], SD1.4 [Rombach et al., 2022b], SD1.5 [Rombach et al., 2022b], SDXL [Rombach et al., 2022b], SSD1B [Gupta et al., 2024], LCM Pixart [Luo et al., 2023], LCM SD1.5 [Luo et al., 2023], LCM SDXL [Luo et al., 2023], SDXL Turbo Sauer et al. [2023] DALLE2 [Ramesh et al., 2022], DALLE3 [Ramesh et al., 2022], IF [DeepFloyd, 2023], Midjourney v5.2 Holz [2023]. Finally, a total of 15$\times$2,000 =30,000 AIGIs are collected. To guarantee diversity, we randomly select 2,000 AIGIs, choosing one AIGI per prompt. Subsequently, we conduct a manual review of these AIGIs to remove any that failed to generate correctly or are unsuitable for annotation. This process results in the final set of AIGIs for A-BenchP1, which can be overviewed in Fig. 6.
@@
A-BenchP2 is designed for the quality evaluation of AIGIs. Consequently, it is essential to ensure that the collected AIGIs span a wide quality range to address various practical scenarios. For Technical Quality, we sample 500 AIGIs from the AIGIQA-20K dataset [Li et al., 2024] using a uniform sampling strategy. Specifically, each AIGI in the AIGIQA-20K dataset is assigned a mean opinion score (MOS) for technical quality. We apply uniform sampling to create more even distributions, as illustrated in Fig. 7. For Aesthetic Quality, in the absence of provided aesthetic scores, we utilize q-align [Wu et al., 2023c], an effective quality predictor, to infer the aesthetic values of AIGIs. Subsequently, we perform uniform sampling similarly to obtain 500 AIGIs for aesthetic evaluation. For Generative Distortion, we manually select 500 AIGIs exhibiting unexpected AIGI-specific distortions. It is important to note that there is no content overlap among the selected AIGIs, which can be overviewed in Fig. 8.
@@ A.4 GPT Evaluation for Choice Judgment
For some LMMs, the response to the question inquiry may vary. For example, given the correct answer C. Blurry’ to the question What is the most severe technical distortion of this image?’, LMMs may respond in different formats: The image is blurry’, There is blur in this image’, or ‘low clarity’. To address the impact of such variations on our evaluation, we’ve implemented a 5-round voting strategy [Wu et al., 2023a]. Under this strategy, we pose the same prompt, as defined in the templates, five times and determine the final outcome based on the majority of GPT’s responses.
@@
[“Score: 1",“Score: 1",“Score: 1",“Score: 1",“Score: 1"]
@@ A.7 Question Type Performance
A.8 Data Statement
@@
A.9 Limitations and Social Impact

### v2

@@ 4.3 Findings of A-Bench
MiniCPM-V2.6 (Qwen2-7B) 69.10% 60.14% 60.47% 64.01%
InternVL2-40B (Nous-Hermes-2-Yi-34B) 66.28% 63.21% 50.10% 59.22%
Ovis1.5 (Llama3-8B) 70.83% 67.82% 55.39% 64.50%
LLaVA-OneVision (Qwen2-7B) 68.84% 67.79% 54.27% 63.78%
CogVLM2-19B (Llama3-8B) 64.21% 61.33% 56.75% 60.73%
IDEFICS-2 (Mistral-7B-Instruct-v0.2) 62.00% 68.76% 47.12% 59.11%
DeepSeek-VL-7B 55.91% 53.79% 47.59% 52.36%
InternLM-XComposer2-VL (InternLM2) 62.29% 63.37% 50.26% 58.58%
LLaVA-NeXT (Llama3-8B) 58.59% 48.57% 52.00% 53.13%
LLaVA-NeXT (Qwen-72B) 59.91% 55.51% 59.80% 58.42%
LLaVA-NeXT (Qwen-110B) 64.69% 57.20% 63.64% 61.89%
mPLUG-Owl2 (LLaMA-7B) 57.90% 54.47% 53.81% 55.45%
LLaVA-v1.5 (Vicuna-v1.5-7B) 45.90% 41.33% 54.59% 47.12%
LLaVA-v1.5 (Vicuna-v1.5-13B) 46.08% 41.22% 48.10% 45.54%
CogVLM-17B (Vicuna-v1.5-7B) 54.76% 48.45% 52.47% 51.36%
Qwen-VL (Qwen-7B) 49.46% 34.34% 50.49% 44.99%
BakLLava (Mistral-7B) 47.88% 33.37% 48.46% 43.39%
Fuyu-8B (Persimmon-8B) 44.61% 30.23% 45.65% 40.20%
random guess 31.87% 32.92% 33.14% 32.63%
@@ 5 Conclusion
6 Ethics Statement
This submission complies fully with the ethical guidelines set by ICLR 2025. We follow ICLR’s principles for responsible AI development, ensuring that our research avoids any potential harm, bias, or discrimination. The data utilized in this work is sourced exclusively from publicly available open-source datasets and models. Furthermore, our methods prioritize fairness, accountability, and transparency in the evaluation of AI-generated images.
7 Limitations
Timeliness Concern
Creating a benchmark involves generating images, collecting data, training evaluators, and verifying data quality, making the process both time-consuming and costly. As a result, it is inevitable that AIGI benchmarks may not always keep pace with the latest technologies or models. However, the insights provided by the benchmark in evaluating AIGI remain valuable and offer useful guidance. We are committed to ongoing updates and expansions to ensure the benchmark remains current.
Scale-up Concern
Since the A-Bench dataset is fully manually annotated and requires validation by at least three other humans, the annotation process is both costly and time-consuming. As such, it is quite challenging to scale up.
8 Acknowledgement
This work is supported in part by the National Natural Science Foundation of China (623B2073, 62101326, 62225112, & 62301316).
@@ A.1 AIGIs Collection
    Bag-of-Words ->Number of Objects Counting: Six logs in a woodpile, stacked so tightly that they seem to form a solid block.
@@
Afterward, we use the collected prompts to create AIGIs. 15 text-to-image generation models are selected, which include: Dreamlike (dreamlike art, 2023), Pixart $\alpha$ Chen et al. (2023a), Playground v2 PlaygroundAI (2023), SD1.4 (Rombach et al., 2022a), SD1.5 (Rombach et al., 2022a), SDXL (Rombach et al., 2022a), SSD1B (Gupta et al., 2024), LCM Pixart (Luo et al., 2023), LCM SD1.5 (Luo et al., 2023), LCM SDXL (Luo et al., 2023), SDXL Turbo Sauer et al. (2023) DALLE2 (Ramesh et al., 2022), DALLE3 (Ramesh et al., 2022), IF (DeepFloyd, 2023), Midjourney v5.2 Holz (2023). Finally, a total of 15$\times$2,000 =30,000 AIGIs are collected. To guarantee diversity, we randomly select 2,000 AIGIs, choosing one AIGI per prompt. Subsequently, we conduct a manual review of these AIGIs to remove any that failed to generate correctly or are unsuitable for annotation. This process results in the final set of AIGIs for A-BenchP1, which can be overviewed in Fig. 6.
@@
A-BenchP2 is designed for the quality evaluation of AIGIs. Consequently, it is essential to ensure that the collected AIGIs span a wide quality range to address various practical scenarios. For Technical Quality, we sample 500 AIGIs from the AIGIQA-20K dataset (Li et al., 2024) using a uniform sampling strategy. Specifically, each AIGI in the AIGIQA-20K dataset is assigned a mean opinion score (MOS) for technical quality. We apply uniform sampling to create more even distributions, as illustrated in Fig. 7. For Aesthetic Quality, in the absence of provided aesthetic scores, we utilize q-align (Wu et al., 2023c), an effective quality predictor, to infer the aesthetic values of AIGIs. Subsequently, we perform uniform sampling similarly to obtain 500 AIGIs for aesthetic evaluation. For Generative Distortion, we manually select 500 AIGIs exhibiting unexpected AIGI-specific distortions. It is important to note that there is no content overlap among the selected AIGIs, which can be overviewed in Fig. 8.
@@ A.4 GPT Evaluation for Choice Judgment
For some LMMs, the response to the question inquiry may vary. For example, given the correct answer C. Blurry’ to the question What is the most severe technical distortion of this image?’, LMMs may respond in different formats: The image is blurry’, There is blur in this image’, or ‘low clarity’. To address the impact of such variations on our evaluation, we’ve implemented a 5-round voting strategy (Wu et al., 2023a). Under this strategy, we pose the same prompt, as defined in the templates, five times and determine the final outcome based on the majority of GPT’s responses.
@@
[“Score: 1”,“Score: 1”,“Score: 1”,“Score: 1”,“Score: 1”]
@@ A.7 Question Type Performance
A.8 Response Variance for LMMs
Considering that the accuracy and stability of the benchmark directly affect the quality of the evaluation, therefore we conduct the response variance experiment here. First, we use a consistent prompt instruction format to minimize any misunderstanding by LMMs and standardize the output. Additionally, we set the model’s temperature parameter to 0, meaning the LMM’s output will no longer be affected by randomness. As a result, the model will give the same response to the same question each time, eliminating variance.
It’s also worth noting that increasing the model’s temperature to encourage more diverse and exploratory answers is indeed an interesting consideration. To further address the concern about the statistical significance of the experiment, we repeat the A-Bench experiment for 5 rounds with different temperature settings across several popular 7B-8B LMMs. The performance is listed in the table below, with the results presented as the mean accuracy ± standard error.
Temperature DeepSeek-VL-7B LLaVA-NeXT-8B LLaVA-v1.5-7B Qwen-VL-7B
0.0 66.58±0.00 67.75±0.00 62.97±0.00 60.41±0.00
0.5 65.11±1.72 66.43±2.09 60.61±2.23 58.17±1.89
1.0 62.04±4.51 63.77±3.86 59.22±4.01 55.22±6.04
Table 4: Performance comparison at different temperatures for various LMM models.
Based on the results, we can observe that when the temperature is set to zero, the accuracy results for all LMMs remain consistent across all 5 rounds. As the temperature increases, the average performance declines and the results become more unstable, with higher standard errors. Therefore, to ensure reproducibility and performance stability, we prefer the zero-temperature setting, as it more accurately and reliably reflects the performance of LMMs, making it more suitable for practical applications.
A.9 Data Statement
@@
A.10 Further Limitations and Social Impact